OpenStudy (anonymous):

A multiple choice test contains 10 questions. There are 4 possible answers for each question. How many ways can a student complete the test if EVERY question must be answered? How many ways can a student complete the test if questions can be left unanswered?

OpenStudy (anonymous):

There are 5 ways to answer each question ( choice 1, 2, 3 , 4 or blank) There are 10 questions.

OpenStudy (anonymous):

So it would be 5*5*5*5*5*5*5*5*5*5=5^10 = 9,765,625